OpenStudy (aaronandyson):

three positive numbers are in the ratio 1/2:1.3:1/5.Find the number id the square of their sum is 244 @sepeario

OpenStudy (sepeario):

If the square of their sum is 244, then what would the sum be?

OpenStudy (sepeario):

Are you still there?

OpenStudy (aaronandyson):

do i add em?

OpenStudy (aaronandyson):

if yes then how?

OpenStudy (aaronandyson):

@Sepeario

OpenStudy (aaronandyson):

u der @Sepeario

OpenStudy (sepeario):

If the square of the sum is 244, then the sum must be root 244, or 15.6. The numbers are in the ratio 0.5:1.3:0.2, yes?

OpenStudy (aaronandyson):

1/2:1/3:1/4

OpenStudy (sepeario):

If the total portion is 2 (0.5+1.3+0.2), then we must divide 15.6 by 2 = 7.8, and to find any of the values, multiply by the selected ratio, e.g for the first number 7.8*0.5 = 3.9

OpenStudy (sepeario):

I know you said 1/3, but for the sake of simplicity and rounding, I kept it as 1.3.

OpenStudy (aaronandyson):

Explain in detail

OpenStudy (aaronandyson):

@Sepeario

OpenStudy (sepeario):

put it like this. \[0.5x+1.3x+0.2x=\sqrt{244}\]

OpenStudy (sepeario):

Solve for x and then you can find the three numbers.

OpenStudy (sepeario):

Can you solve from there?

OpenStudy (aaronandyson):

its1/3 not 1.3 my typos X_X

OpenStudy (sepeario):

Yeah but when you round it it is 1.3 :)

OpenStudy (sepeario):

oh wait no it's 0.3 oops

OpenStudy (sepeario):

So yeah can you solve the equation though?

OpenStudy (aaronandyson):

1x = sqrt244???

OpenStudy (aaronandyson):

@Sepeario

OpenStudy (sepeario):

good

OpenStudy (sepeario):

then x =?

OpenStudy (aaronandyson):

15.62?

OpenStudy (sepeario):

cool then to find each value multiply by the coefficient i.e. 0.2, 0.5 etc.

OpenStudy (aaronandyson):

The back of book says its 12,8 and 6?:/

OpenStudy (sepeario):

i'm sorry but i honestly think that i don't know then.. :(